OpenStudy (anonymous):

Suppose a triangle has two sides of length 42 and 35, and that the angle between these two sides is 120. What is the length of the third side of the triangle? a.66.78 b.58.04 c.43.65 d.77

OpenStudy (campbell_st):

Use the law of cosines \[a^2 = b^2 + c^2 - 2 \times b \times c \times \cos(A)\] in your question b = 42, c = 35 and A = 120 just substitute and calculate

OpenStudy (anonymous):

thank you i got it!